Suppose u and v are nonzero vectors in 3-space, where u=(u1, u2, u3) and v=(v1, v2, v3). Prove that u x v is perpendicular to both u and v by making use of the dot product.
1
Expert's answer
2020-04-06T15:57:40-0400

"c=u\times v=\begin{vmatrix}
 i & j& k \\
 u_1 & u_2 & u_3\\
 v_1 & v_2 & v_3
\end{vmatrix}=(u_2v_3-u_3v_2)i-(u_1v_3-u_3v_1)j+(u_1v_2-u_2v_1)k"

"u\cdot c=u_1(u_2v_3-u_3v_2)-u_2(u_1v_3-u_3v_1)+u_3(u_1v_2-u_2v_1)=0"

Which means u is perpendicular to c.

"v\cdot c=v_1(u_2v_3-u_3v_2)-v_2(u_1v_3-u_3v_1)+v_3(u_1v_2-u_2v_1)=0"

Which means v is perpendicular to c.